OpenStudy (anonymous):

can someone help me factor x^2 + 6x + 72?

jhonyy9 (jhonyy9):

x_1,_2=(-6+/-sqrt(36-288))2=(-6+/-sqrrt(-252))/2=(-6+/- isqrt(4*9*7)/2 = =2(-3+/- 3isqrt7)/2 = -3+/- 3isqrt7

jhonyy9 (jhonyy9):

= -3(1+/- isqrt7)

jhonyy9 (jhonyy9):

than this will be -3(x-(1-isqrt7))(x+(1+isqrt7))